6. Fox Company manufactures a specialty line of belts. The company uses a job-order costing system. Cost detail for the job follow: Units produced 5,000 belts Direct materials in the amount of $43,000 Direct labor in the amount of $18,500 The POHR was $25.00 per machine-hour (MH) and Job A10 required 400 MHs. Selling and G & A expenses of $23,000 were incurred for the month. By the end of the month Fox had sold 4,000 of these belts and overhead was over applied by a total of $5,000. The cost of goods sold for the month, computed using the direct method (including any necessary adjusting entry for over or under applied overhead) was: A) $75,200 B) $71,500 C) $57,200 D) $52,200
Answer #1

Answers

  • Applied overhead = 400 MHs x $ 25 per MH = $ 10,000
  • Overheads are over applied by $ 5000 [mentioned in question]. Hence, actual Overhead = 10000 – 5000 = $ 5,000
  • Cost of Goods manufactured = 43000 + 18500 + 10000 applied overhead = $ 71,500
  • Units produced = 5000 belts
  • Unadjusted Cost of Goods Sold = ($71500 / 5000 belts) x 4000 belts
    = $ 57,200
  • Adjusted Cost of Goods Sold = $ 57200 - $ 5000 over applied overhead = $ 52,200
  • Correct Answer = Option ‘D’ $ 52,200